### v3

# Large deviations of current for the symmetric simple exclusion process on a semi-infinite line, and on an infinite line with a slow bond

Abstract
Two influential exact results in classical one-dimensional diffusive transport are about current statistics for the symmetric simple exclusion process: one in the stationary state on a finite line coupled with two unequal reservoirs at the boundaries, and the other in the non-stationary state on an infinite line. We present the corresponding result for the intermediate geometry of a semi-infinite line coupled with a single reservoir. This result is obtained using the fluctuating hydrodynamics approach of macroscopic fluctuation theory and confirmed by rare event simulations using a cloning algorithm. We apply our exact result for solving several related challenging problems, namely, the full counting statistics in presence of a defect bond, exclusion process with localized injection, survival of a tagged particle in presence of an absorbing boundary, and the stretched exponential decay in a kinetically constrained model.
large deviations, macroscopic fluctuation theory, symmetric simple exclusion process, slow bond, kinetically constrained model, stretched exponential correlation, target survival, fast injection
preprint: TIFR/TH/24-4
Current fluctuations in non-equilibrium transport have long been a subject of interest in both classical and quantum contexts [2002_Prähofer_Current, 2005_Bertini_Current, *2006_Bertini_Non, 2007_Derrida_Non, 2015_Lazarescu_The, 2020_Banerjee_Current, 2023_Bello_Current, 1995_Lee_Universal, 2023_McCulloch_Full, 2024_Wienand_Emergence, turkeshi2024]. The major interest in these investigations lies in the full counting statistics in terms of large deviations of current. Apart from being a contender for extension of free energy for non-equilibrium systems, large deviation function (ldf) in general can characterize various peculiarities of non-equilibrium conditions, such as non-local response, emergent symmetries, and low-dimensional phase transitions [2005_Bodineau_Distribution, 2007_Derrida_Non, 2008_Bodineau_Long, 2010_Jona_From, 2009_Touchette_The, 2009_Bertini_Towards, 2011_Hurtado_Symmetries, 2018_Baek_Dynamical]. However, estimating ldf poses a major challenge, often requiring specialized integrability techniques [2001_Schutz_Exactly, 2015_Lazarescu_The, 2015_Mallick_The] tailored to curated models or clever numerical sampling schemes of rare events [2006_Giardinà_Direct, 2007_Lecomte_Numerical, 2011_Giardinà_Simulating, 2019_Perez_Sampling, Hartmann_computer_sim]. Understandably, exact results about ldf play an important role in the landscape of non-equilibrium physics, providing a benchmark for affirming qualitative predictions of approximate methods. Our work in this Letter presents a non-trivial addition to this list of exact results, fostering solution to further problems of practical significance [1989_Spohn_Stretched, 2012_Krapivsky_Symmetric, 2013_Franco_Phase, *2016_Franco_Phase].
Among the widely studied stochastic models of classical transport, is the symmetric simple exclusion process (SSEP) [1991_Spohn_Large, 1999_Kipnis_Scaling, 1999_Liggett_Stochastic, 2007_Derrida_Non]. SSEP, along with its driven variants, has attained the status of a paradigmatic model in non-equilibrium statistical physics [2011_Chou_Non, 2015_Mallick_The]. Two celebrated exact results for SSEP concern the large deviations of current on a finite lattice coupled with two unequal reservoirs, and on an infinite lattice starting with a non-stationary state. The two geometries represent distinct non-equilibrium scenarios: a stationary state for finite systems, and a time-dependent state relaxing towards an asymptotic equilibrium state for infinite systems. They illuminate crucial differences in their fluctuations. The finite system, in the long run, holds no memory of the initial state, while the infinite system exhibits an unusual dependence on the initial state even at long times [2009_Derrida_Current2, 2012_Krapivsky_Fluctuations].
The large deviations of current in these two geometries were obtained using the additivity conjecture [2004_Bodineau_Current, 2007_Bodineau_Cumulants, 2021_Derrida_Large] and integrability methods, such as the diagonalization of tilted matrix [2004_Derrida_Current] or the matrix product states [2012_Gorissen_Exact, 2015_Lazarescu_The] for the finite lattice, and the Bethe ansatz [2009_Derrida_Current] for the infinite lattice. These microscopic results were subsequently verified [2005_Bertini_Current, *2006_Bertini_Non, 2022_Mallick_Exact, *2024_Mallick_Exact, 2024_Saha_Large] using a fluctuating hydrodynamics framework [2009_Derrida_Current, 2015_Bertini_Macroscopic, 2023_Jona_Current]. This framework, now famously known as the macroscopic fluctuation theory (MFT) [2001_Bertini_Fluctuations, *2002_Bertini_Macroscopic, 2005_Bertini_Current, *2006_Bertini_Non], emerged in the early 2000s from the seminal works of Bertini, De Sole, Gabrielli, Jona-Lasinio, and Landim, which presented a general approach for characterizing non-equilibrium fluctuations of diffusive systems. MFT has successfully led to exact results of large deviations in exclusion processes [2005_Bertini_Current, *2006_Bertini_Non, 2007_Tailleur_Mapping, *2008_Tailleur_Mapping, 2015_Bertini_Macroscopic, 2022_Mallick_Exact, 2024_Saha_Large] and related transport models [2015_Bertini_Macroscopic, 2022_Bettelheim_Inverse, *2022_Bettelheim_Full, 2023_Agranov_Macroscopic, 2024_Bettelheim_Complete].
In this Letter, we consider the intermediate scenario: SSEP on a semi-infinite lattice [1975_Liggett_Ergodic, 2004_Grosskinsky_Phase] coupled to a boundary reservoir with a density that is different from the initial bulk density of the system. This elucidates a non-equilibrium regime evolving into an asymptotic state in equilibrium with the reservoir [2023_Saha_Current]. In fact, the current statistics in this geometry describes the early time ($t\ll L^{2}$) statistics near the reservoir of a finite system of length $L$, extending the previous steady state ($t\gg L^{2}$) result [2004_Bodineau_Current]. The statistics have also been useful recently in the context of Mpemba effect [turkeshi2024] in quantum circuits and particle injection in a lattice [2014_Krapivsky_Lattice, 2012_Krapivsky_Symmetric]. We shall further see that the semi-infinite geometry has direct relation to transport in presence of defect and survival statistics of tracer. Despite the relevance, only limited results [2012_Williams_Combinatorics, 2013_Tracy_The, 2018_Duhart_The, 2023_Saha_Current] are known for this intermediate geometry, particularly because extending the aforementioned integrability methods for this geometry proves challenging. Even a solution via MFT has remained elusive [2023_Saha_Current].
We overcame these technical challenges by uncovering a novel mapping to the infinite-line problem that leads to an exact result for the ldf of current for SSEP on the semi-infinite line, which is our main result. The mapping to the infinite-line-problem is possible due to certain symmetries of the Euler-Lagrange equations and the boundary conditions of the associated variational problem within MFT. These techniques, however, could not be obviously exploited within microscopic approaches. Our exact result is a testament to the power of MFT for non-equilibrium systems that are otherwise formidable to approach.
Our exact result for the semi-infinite SSEP enables the solution of three additional challenging problems. First, we obtain the exact ldf of current for an infinite SSEP with a localized slow bond, extending the celebrated result of Derrida and Gerschenfeld [2009_Derrida_Current]. Second, we derive the corresponding ldf for fast particle injection at a single site [2014_Krapivsky_Lattice, 2012_Krapivsky_Symmetric] in an infinite SSEP. Third, we compute the survival probability of a tracer in a fully packed SSEP with an absorbing boundary. This result extends the classical static-target [2014_Meerson_Survival] survival problem, which concerns the special case where no particle escapes the system.
Model and main result: SSEP on a semi-infinite lattice (see Fig. 1), with sites indexed by $i=1,2,\cdots$, is composed of continuous-time hard-core random walkers hopping to adjacent sites with unit rate provided the target site is empty. At the boundary site $i=1$, particles are injected following exclusion with rate $\gamma\rho_{a}$ and removed at rate $\gamma(1-\rho_{a})$, which models [2007_Derrida_Non] coupling with a reservoir of density $\rho_{a}$. For each site $i$ at a given time $\tau$, we assign a binary occupation number $n_{i}(\tau)$ that takes values 0 and $1$, depending on whether the site is empty or occupied, respectively. Initially, sites are filled following Bernoulli distribution with a uniform average density $\rho_{b}$.
Figure 1: SSEP on a semi-infinite lattice coupled with a reservoir of density $\rho_{a}$ with a coupling strength $\gamma>0$.
Our main result concerns the time-integrated current, $Q_{T}$, which represents the total flux of particles from the reservoir into the system over a time-period $T$. In the hydrodynamic description [SM], expressed in terms of coarse-grained density $\rho(i/\sqrt{T},\tau/T)\simeq n_{i}(\tau)$, the flux
$Q_{T}={\sqrt{T}}\int_{0}^{\infty}\mathrm{d}x\,[\rho(x,1)-\rho(x,0)]$ (1)
measures the net change in the number of particles in the system. In the large $T$ limit, its generating function has the asymptotics
$\big<\mathrm{e}^{\lambda Q_{T}}\big>\sim\mathrm{e}^{\sqrt{T}\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})}.$ (2)
The scaled cumulant generating function (scgf)
$\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})=R_{\text{si}}\big(\omega(\lambda,\rho_{a},\rho_{b})\big)$ (3a)
(subscript ‘si’ denotes semi-infinite) with
$R_{\text{si}}(\omega)=\begin{cases}R(\omega)&\text{for }\omega\geq-\frac{1}{2}\\ \displaystyle{-R(\omega)-\frac{1}{\sqrt{\pi}}\zeta\bigg(\frac{3}{2}\bigg)}&\text{for }\omega\leq-\frac{1}{2}\end{cases}$ (3b)
where $\zeta(z)$ is the Riemann zeta function and
$R\big(\omega\big)=\int_{-\infty}^{\infty}\frac{\mathrm{d}k}{2\pi}\log{\big[1+4\,\omega\,(1+\omega)\,\mathrm{e}^{-k^{2}}\big]}$ (3c)
and $\omega$ is a function of the parameters defined as [2009_Derrida_Current, 2012_Krapivsky_Fluctuations]
$\omega(\lambda,\rho_{a},\rho_{b})=\rho_{a}(1-\rho_{b})(\mathrm{e}^{\lambda}-1)+\rho_{b}(1-\rho_{a})(\mathrm{e}^{-\lambda}-1).$ (3d)
This parametric dependence of scgf on $\lambda$, $\rho_{a}$, and $\rho_{b}$ through a single function $\omega$ arises from a symmetry of the underlying dynamics of SSEP [2004_Derrida_Current, 2009_Derrida_Current, 2010_Lecomte_Current]. The piecewise function (3b) is constructed by analytic continuation of $R(\omega)$ demanding a convex scgf independently confirmed by numerical simulation (see End Matter). The expression (3) is consistent with an earlier result [2023_Saha_Current] in the low density limit.
The function $R(\omega)$ has different representations. For example, $R(\omega)=-1/(2\sqrt{\pi})\mathrm{Li}_{3/2}(-4\omega(1+\omega))$ in terms of the poly-logarithm function $\mathrm{Li}_{s}(z)$ is similar to the expression of the number of particles in Fermi-Dirac statistics. Another representation in terms of Fredholm determinant
$R(\omega)=\frac{1}{2\pi}\log{\det{[\mathbbm{1}+4\omega(1+\omega)K(x,y)]}},$ (4)
with the kernel $K(x,y)=\delta(x-y)\exp{(\partial_{x}^{2})}$, finds analogue in the current statistics for asymmetric exclusion process [2000_Johansson_Shape, 2008_Tracy_A, 2009_Tracy_Total].
It is instructive to compare (3) with the corresponding result for the infinite line [2009_Derrida_Current]. Similar to the infinite line case [2009_Derrida_Current2], the scgf (3) admits the Gallavotti-Cohen-type fluctuation symmetry [Evans1993PRL, Gallavotti1995PRL, Kurchan1998JPA, Lebowitz1999JSP]
$\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})=\mu_{\text{si}}\bigg(\log{\frac{\rho_{b}\,(1-\rho_{a})}{\rho_{a}\,(1-\rho_{b})}}-\lambda,\rho_{a},\rho_{b}\bigg).$ (5)
Figure 2: Scgf for semi-infinite SSEP: Solid red line represents the theoretical result of scgf (3) for $\rho_{a(b)}=0.5$, with blue dots representing the corresponding simulation result obtained by the cloning algorithm. The results closely match in the range $1\leq\lambda\leq 1$, with the deviations shown in the inset. For comparison, Gaussian approximation and the scgf for infinite line $\mu_{\text{inf}}$ [2009_Derrida_Current] are shown in dotted magenta and black dashed line, respectively.
Numerical confirmation: For confirming our result (3) based on the hydrodynamic theory for SSEP, we have independently generated the scgf using rare-event simulation based on a continuous-time cloning algorithm [SM, 2006_Giardinà_Direct, 2007_Lecomte_Numerical] with $10^{5}$ clones and measured over duration $T=250$. The data, plotted in Fig. 2, shows a good agreement with our theoretical result (3) for a reasonably large parameter range of $\lambda$. Deviations emerging at larger values of $\lambda$ are a consequence of finite-size effects [SM].
LDF: The scaling in (2) corresponds to a large deviation asymptotics of the probability
$P\big(Q_{T}=j\,\sqrt{T}\big)\sim\mathrm{e}^{-\sqrt{T}\phi_{\text{si}}(j)}$ (6)
where $\phi_{\text{si}}(j)$ is the ldf, related to $\mu_{\text{si}}(\lambda)$ (reference to $\rho_{a(b)}$ is ignored) by a Legendre-Fenchel transformation [2009_Touchette_The, SM] $\phi_{\text{si}}(j)=\max_{\lambda}{[j\lambda-\mu_{\text{si}}(\lambda)]}$. Using the transformation, it is immediate that (5) reflects the symmetry
$\phi_{\text{si}}(j)-\phi_{\text{si}}(-j)=j\log{\frac{\rho_{b}\,(1-\rho_{a})}{\rho_{a}\,(1-\rho_{b})}}.$ (7)
Similarly, the asymptotics $\mu_{\text{si}}(\lambda)\simeq[(4\sqrt{2})/(3\pi)](\log{\omega})^{3/2}$ of (3) for large positive $\lambda$ correspond [2009_Derrida_Current] to the asymptotics of the ldf
$\phi_{\text{si}}(j)\simeq\frac{\pi^{2}j^{3}}{24}-j\log{\big(\rho_{a}\,(1-\rho_{b})\big)}$ (8)
for large positive $j$. A similar analysis for large negative $j$ gives asymptotics (8) with $j\to-j$ and $\rho_{a}\leftrightarrow\rho_{b}$. Notably, for the non-equilibrium condition $\rho_{a}\neq\rho_{b}$, the ldf is skewed.
Derivation: In the following, we outline our derivation of (3) within the fluctuating hydrodynamics framework of MFT. The crucial idea behind MFT is to recognize the relevant hydrodynamic modes for a coarse-grained description of the dynamics and characterize the probability of their evolution in terms of an Action, which is analogous to the Martin-Siggia-Rose-Janssen-De Dominicis (MSRJD) Action [1973_Martin_Statistical, 1976_Janssen_On, 1978_Dominicis_Dynamics, 1978_Dominicis_Field] of the associated fluctuating hydrodynamics equation. For SSEP, the relevant hydrodynamic mode is the locally conserved density $\rho(x,t)$ evolving by [2023_Saha_Current, SM]
$\partial_{t}\rho=\partial_{x}^{2}\rho+\frac{1}{T^{1/4}}\,\partial_{x}\big(\sqrt{\sigma(\rho)}\,\eta\big)$ (9)
where $\sigma(\rho)=2\rho(1-\rho)$, and $\eta(x,t)$ is a delta-correlated Gaussian white noise with unit covariance. Corresponding MSRJD-Action on the semi-infinite line is [2023_Saha_Current, SM]
$S_{0}=\mathcal{F}+\int_{0}^{1}\mathrm{d}t\int_{0}^{\infty}\mathrm{d}x\bigg[\widehat{\rho}\,\partial_{t}\rho-\bigg(\frac{\sigma(\rho)}{2}\partial_{x}\widehat{\rho}+\partial_{x}\rho\bigg)\partial_{x}\widehat{\rho}\bigg$] (10)
where $\widehat{\rho}$ is the response field and $\mathcal{F}$ incorporates contributions from fluctuations in the initial state [2009_Derrida_Current2, 2023_Saha_Current]. Within this description, the generating function $\langle\mathrm{e}^{\lambda Q_{T}}\rangle=\int\mathcal{D}\left[\rho,\widehat{\rho}\right]\,\mathrm{e}^{-\sqrt{T}S_{\lambda}\left[\rho,\widehat{\rho}\right]}$ with $S_{\lambda}=S_{0}-\lambda Q_{T}/\sqrt{T}$ and $Q_{T}$ in (1).
For large $T$, the path-integral is dominated by a saddle point, leading to (2) with
$\mu_{\text{si}}(\lambda)\simeq-\min_{\rho,\widehat{\rho}}{S_{\lambda}\left[\rho,\widehat{\rho}\right]}\equiv-S_{\lambda}\left[q_{\text{si}},p_{\text{si}}\right]$ (11)
[…]
$\partial_{t}q_{\text{si}}$ $=\partial_{x}^{2}q_{\text{si}}-\partial_{x}[\sigma(q_{\text{si}})\,\partial_{x}p_{\text{si}}],\quad\text{and}$ (12a)
$\partial_{t}p_{\text{si}}$ $=-\partial_{x}^{2}p_{\text{si}}-\frac{\sigma^{\prime}(q_{\text{si}})}{2}\,(\partial_{x}p_{\text{si}})^{2}$ (12b)
in the semi-infinite domain $x>0$, with the temporal boundary conditions
$p_{\text{si}}(x,0)=\lambda+\int_{\rho_{b}}^{q_{\text{si}}(x,0)}\frac{2\,\mathrm{d}r}{\sigma(r)}\quad\text{and}\quad p_{\text{si}}(x,1)=\lambda,$ (13)
where the integral in $p_{\text{si}}(x,0)$ is the contribution [2009_Derrida_Current2] from $\mathcal{F}$ for the initial state of SSEP with Bernoulli measure. Additional spatial boundary conditions,
$q_{\text{si}}(0,t)=\rho_{a}\quad\text{and}\quad p_{\text{si}}(0,t)=0$ (14)
are due to the “fast-coupling” [2024_Saha_Large, SM] with reservoir.
This variational problem is reminiscent of the corresponding problem on the infinite line [2009_Derrida_Current2] for a Bernoulli-measured initial state with an average density $\rho_{a}$ for $x<0$ and $\rho_{b}$ for $x>0$. The only differences between the two problems are the domain $x$ and the boundary conditions. The infinite line problem was formulated in [2009_Derrida_Current2] within MFT, which was recently solved in [2022_Mallick_Exact] by identifying an ingenious mapping to the classical integrable system and employing the inverse scattering method. Despite having small differences, the semi-infinite case poses a new nontrivial problem [1975_Ablowitz_The], which is incredibly difficult to solve.
A simplification arises for the special choice of initial density pair $(\rho_{a},\rho_{b})\equiv(1/2,0)$ for the semi-infinite line problem. For this choice of densities, there are no fluctuations in the initial state, which amounts to setting $\mathcal{F}=0$ in (10) with the condition
$q_{\text{si}}(x,0)=0.$ (15)
The Euler-Lagrange equations for the corresponding variational problem (11) remains the same as in (12). The only difference comes in the initial condition in (13), which is now replaced by (15). This initial condition corresponds to the quenched averaging [2009_Derrida_Current2, 2024_Saha_Large].
We now show that this quenched semi-infinite line problem has a direct mapping to the quenched infinite line problem with $(\rho_{a},\rho_{b})\equiv(1,0)$ and fugacity $\tilde{\lambda}=2\lambda$. For the latter problem, the Euler-Lagrange equation is the same [2009_Derrida_Current2] as in (12), but now on the entire real line $x$ with the temporal boundary conditions
$q_{\text{inf}}(x,0)=\theta(-x)\quad\text{and}\quad p_{\text{inf}}(x,1)=\tilde{\lambda}\,\theta(x)$ (16)
(subscript ‘inf’ denotes infinite and $\theta(x)$ is the Heaviside step function). It is simple to check that the solution admits the symmetry
$q_{\text{inf}}(x,t)$ $=1-q_{\text{inf}}(-x,t)$ (17a)
$p_{\text{inf}}(x,t)$ $=\tilde{\lambda}-p_{\text{inf}}(-x,t)$ (17b)
which fixes the value of the fields at the origin
$q_{\text{inf}}(0,t)=\frac{1}{2}\quad\text{and}\quad p_{\text{inf}}(0,t)=\frac{\tilde{\lambda}}{2}$ (18)
at all times $0<t<T$. This conclusion is an essential part of our observation, as now, the solution $q_{\text{inf}}(x,t)$ on positive line $x$ satisfies the boundary conditions (14, 15) of the semi-infinite line problem with $(\rho_{a},\rho_{b})\equiv(1/2,0)$. For a similar correspondence of the response field, we define
$\tilde{p}(x,t)=p_{\text{inf}}(x,t)-\frac{\tilde{\lambda}}{2}$ (19)
[…]
Consequently, the least-Action path $(q_{\text{si}},p_{\text{si}})$ for the semi-infinite line problem with $(\rho_{a},\rho_{b})\equiv(1/2,0)$ and fugacity $\lambda$ is related to the corresponding path $(q_{\text{inf}},p_{\text{inf}})$ for the infinite line problem with $(\rho_{a},\rho_{b})\equiv(1,0)$ and fugacity $2\lambda$ by
$q_{\text{si}}(x,t)=q_{\text{inf}}(x,t)\quad\text{and}\quad p_{\text{si}}(x,t)=p_{\text{inf}}(x,t)-\lambda$ (20)
for $x\geq 0$ at all times. This correspondence relates the least-Action of the two problems resulting [SM] in our crucial observation:
$\mu_{\text{si}}\Big(\lambda,\frac{1}{2},0\Big)=\frac{1}{2}\,\mu_{\text{inf}}(2\lambda,1,0)$ (21)
where the pre-factor $1/2$ comes from the half-domain of integration of $x$ in (10). The latter scgf is known from the seminal work [2009_Derrida_Current] of Derrida and Gerschenfeld, which culminates in
$\mu_{\text{si}}\Big(\lambda,\frac{1}{2},0\Big)=\int_{-\infty}^{\infty}\frac{\mathrm{d}k}{2\pi}\,\log{[1+(\mathrm{e}^{2\lambda}-1)\mathrm{e}^{-k^{2}}]}.$ (22)
The result (22) is for the specific initial density pair $(\rho_{a},\rho_{b})\equiv(1/2,0)$. For extending the result for other densities, we invoke a well-known rotational symmetry [2009_Derrida_Current2, 2010_Lecomte_Current, SM] of the least-Action (11). Essentially, the least-Action paths for two sets of parameters $(\lambda,\rho_{a},\rho_{b})$ and $(\lambda^{\prime},{\rho_{a}}^{\prime},{\rho_{b}}^{\prime})$ are related under a canonical transformation [2009_Derrida_Current2, 2010_Lecomte_Current, SM]. The symmetry results [2009_Derrida_Current, 2010_Lecomte_Current, SM] in a dependence (3a) of the scgf on $(\lambda,\rho_{a},\rho_{b})$ through a single parameter $\omega$ defined in (3d). This dependence enables us [SM] to deduce the function $R_{\text{si}}(\omega)$ from the result (22), leading to the expression (3b). Note that the $\omega$-dependence does not extend [2004_Derrida_Current, 2022_Mallick_Exact] for optimal profile $q_{\textrm{si}}(x,t)$, making it difficult to infer the profile at arbitrary densities.
Slow bond: Recent interests [2017_Baldasso_Exclusion, 2019_Franco_Non, 2020_Goncalves_Non, 2020_Erignoux_HydrodynamicsI, *2020_Erignoux_HydrodynamicsII, 2021_Derrida_Large, 2024_Saha_Large, 2023_Saha_Current] in studying the effects of the coupling strength with reservoir can also be addressed in the semi-infinite line problem. The result (3) is independent [2023_Saha_Current] of the coupling strength $\gamma$ (see Fig. 1) as long as it is larger than $\mathrm{O}(T^{-1/2})$. This is seen from the corresponding result for slow coupling $\gamma=\Gamma/\sqrt{T}$, where the boundary-fluctuations are significant, modifying the scgf (3a) to
$\mu_{\text{si}}^{\text{slow}}(\lambda,\rho_{a},\rho_{b})=\min_{z}{\Big[\Gamma\sinh^{2}{(z-u)}+R_{\text{si}}(\sinh^{2}{z})\Big]}$ (23)
with $\sinh^{2}{u}=\omega(\lambda,\rho_{a},\rho_{b})$. In the $\Gamma\to\infty$ limit, (3a) is recovered from (23). The expression (23) is obtained following an additivity argument [2021_Derrida_Large], where contributions in (23) are separately from the single bond joining the reservoir and the system, and the system itself, optimised over the density at their common site [SM]. This construction is very similar to the discussion in [2021_Derrida_Large] for a related context, except for a crucial distinction that unlike the latter example, there is no quasi-stationarity for the semi-infinite problem. Despite this, the additivity conjecture gives the correct result (23).
A similar additivity argument helps to solve the problem of current fluctuation across a single slow bond at the origin on an infinite one-dimensional lattice [2013_Franco_Hydrodynamical, 2013_Franco_Phase, *2016_Franco_Phase, 2020_Erhard_Non], initially filled with uniform average density $\rho_{a(b)}$ in the negative (positive) half-line. For the hopping rate $\gamma=\Gamma/\sqrt{T}$ across the slow bond, and unity for rest of the lattice, the generating function of current $\big<\mathrm{e}^{\lambda Q_{T}}\big>\sim\mathrm{e}^{\sqrt{T}\mu_{\text{inf}}^{\text{slow}}(\lambda,\rho_{a},\rho_{b})}$, for large $T$, with the scgf [SM]
$\mu_{\text{inf}}^{\text{slow}}$ $(\lambda,\rho_{a},\rho_{b})=\min_{z_{a},z_{b}}\Big[R_{\text{si}}\big(\sinh^{2}{z_{a}}\big)$
$+\Gamma\sinh^{2}{(z_{a}+z_{b}-u)}+R_{\text{si}}\big(\sinh^{2}{z_{b}}\big)\Big]$ (24)
where $\sinh^{2}{u}=\omega(\lambda,\rho_{a},\rho_{b})$. In the $\Gamma\to\infty$ limit, the celebrated infinite line result [2009_Derrida_Current, 2022_Mallick_Exact] is recovered [SM] from (24).
Other applications: Several other interesting conclusions can be drawn from our result. A known symmetry [2009_Derrida_Current2] of the MFT-Action extends our exact results to models [1982_Kipnis_Heat, 2007_Giardinà_Duality, 2014_Vafayi_Weakly, 2022_Franceschini_Symmetric] with a quadratic mobility $\sigma(\rho)=2A\rho(B-\rho)$, where $A$ and $B$ are arbitrary constants, culminating in $\mu_{\sigma}(\lambda,\rho_{a},\rho_{b})=(1/A)\,\mu_{\text{ssep}}\big(\omega(AB\lambda,\rho_{a}/B,\rho_{b}/B)\big)$. Appropriate limits of $A$ and $B$ give the result for the Kipnis-Marchioro-Presutti (KMP) model [1982_Kipnis_Heat, 2007_Giardinà_Duality], the symmetric simple inclusion process (SSIP) [2014_Vafayi_Weakly, 2022_Franceschini_Symmetric], and the non-interacting particles [2023_Saha_Current].
The semi-infinite line result provides solutions for two closely related problems. The first problem [2012_Krapivsky_Symmetric, 2014_Krapivsky_Lattice] concerns an infinite line SSEP with fast injection at a single site, a scenario which serves as the simplest description of hard-core particles spreading on an initially empty system [2019_Krapivsky_Free, *2020_Krapivsky_Free]. Such dynamics arise in diverse contexts, including the voter model [2003_Mobilia_vote], chemical catalysis [1992_Krapivsky_monomer, 1996_Krapivsky_catalytic], and wetting thin film spreading [Popescu_2012_wetting_films]. At long-times $t$, the net injection of particles $N_{t}$ on an empty lattice follows the asymptotics $\langle\mathrm{e}^{\lambda N_{t}}\rangle\sim\exp{[2\sqrt{t}\mu_{\text{si}}(\lambda,1,0)]}$ with (3), which results from the statistical independence of the two halves due to fast injection. The second problem concerns the survival probability $P_{t}(k)$ of the $k^{\text{th}}$ tagged particle (tracer) up to time $t$ in a fully packed SSEP on $\mathbb{Z}^{+}$ in the presence of an absorbing site at the origin. The problem has applications to first passage [2013_Majumdar_first_passage, 2014_Benichou_first_passage] problems, target search [2011_Benichou_search], and diffusion limited reactions [rice1985diffusion]. The asymptotics (6) and the relation $P_{t}(k)=P(-Q_{t}>k)$ in the semi-infinite SSEP gives a stretched exponential decay $P_{t}(k)\sim\exp{(-\sqrt{t/\tau_{s}})}$ for $k\lesssim 2\sqrt{t/\pi}$ at long-times $t$, with $\tau_{s}=[\phi(-k/\sqrt{t})]^{-2}$.
The final application we discuss is about similar stretched exponential decays in kinetically constrained models [1983_Skinner_Kinetic, 1989_Spohn_Stretched, 2020_Gupta_Size, 2024_Mukherjee_Stretched], such as the energy-conserving spin-flip dynamics [1989_Spohn_Stretched, 2024_Mukherjee_Stretched], where spin auto-correlation decays as $\exp(-\sqrt{t/\tau})$, with $\tau$ remained undetermined. For the spin-model, rigorous result about $\tau$ comes [SM] from realizing that the domain-wall dynamics is equivalent to the infinite-line SSEP at a uniform density $\rho$. This correspondence relates [SM] the spin-auto-correlation to the generating function of current in the infinite SSEP leading to $\tau=[\mu_{\text{inf}}(\mathrm{i}\pi,\rho,\rho)]^{-2}$. Similar correspondence holds in presence of a slow bond or in semi-infinite geometry.
Conclusions: We derived full counting statistics (3) for SSEP on a semi-infinite line, which represents an intermediate non-equilibrium regime between the well-studied finite [2004_Bodineau_Current, 2005_Bertini_Current] and infinite SSEP [2009_Derrida_Current, 2022_Mallick_Exact]. Compared to the latter examples, the scgf (3) exhibits a piecewise solution (3b). Our exact result not only solves the absorbing boundary problem relevant to target survival [2011_Benichou_search], but also leads to the full counting statistics (24) for the infinite-SSEP with a defect bond. The variational construction used for the latter solution offers a powerful approach for studying inhomogeneous transport, even for asymmetric dynamics.
There remain several related open problems of immediate interest. Most prominent among them is the scgf for fixed (quenched) initial states. The only available non-trivial result [2009_Derrida_Current2] is for the infinite line SSEP at half-filling, obtained following a relation with the fluctuating (annealed) initial state. A similar mapping for the semi-infinite SSEP yields the quenched scgf
$\mu_{\text{si}}^{\text{que}}\Big(\lambda,\frac{1}{2},\frac{1}{2}\Big)=\int_{-\infty}^{\infty}\frac{\mathrm{d}k}{\sqrt{8}\,\pi}\log{\big[1+\sinh^{2}{(\lambda)}\,\mathrm{e}^{-k^{2}}\big]}.$ (25)
An extension of these results in higher dimensions would be interesting, where only limited results are available [2013_Akkermans_Universal, Becker_2015, Hutardo_2016, Hutardo_2017, Harris_2016]. From a practical point of view, the emergence of the SSEP in quantum circuits [2023_McCulloch_Full] or hydrodynamics in chaotic quantum systems [2024_Wienand_Emergence] offers exciting applications, where semi-infinite analog [turkeshi2024] becomes relevant. Along similar lines, extensions of our results for quantum analogues of SSEP [2019_Bernard_Open, 2021_Bernard_Can] or for integrable models [2020_Doyon_Lectures, 2022_Bastianello_Introduction] would be timely.
Acknowledgements.
Acknowledgments: We acknowledge the financial support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI 4002. TS thanks Bernard Derrida for insightful discussions on the problem. The importance of semi-infinite line SSEP and the additivity ansatz for slow bond originated from those discussions. TS also thanks Mustansir Barma for introducing him to the works on stretched exponential decays, particularly the open problem in [1989_Spohn_Stretched]. TS thanks the support from the International Research Project (IRP) titled ‘Classical and quantum dynamics in out of equilibrium systems’ by CNRS, France.
I End Matter
The piecewise function: The scgf (3) is plotted in Fig. 3 for parameter regimes where the piecewise nature of the function (3b) is relevant. The function $R(\omega)$ in (3c) is singular at $\omega=-1/2$, and this leads to a non-convex $R(\omega(\lambda,\rho_{a},\rho_{b}))$ as a function of $\lambda$, indicated by the dashed line in Fig. 3. The piecewise function (3b) constructed from $R(\omega)$ gives an analytic, convex scgf, which agrees with the data from rare event simulation using a cloning algorithm.
Figure 3: Scgf for semi-infinite SSEP: The scgf of current for the semi-infinite SSEP for densities $(\rho_{a},\rho_{b})\equiv(0.9,0.1)$. The solid red line represents the theoretical result (3), while the data points are obtained by the cloning algorithm with $T=500$ and $N_{c}=50,000$. The dashed line represents $R(\omega(\lambda,\rho_{a},\rho_{b}))$ in (3c) as a function of $\lambda$ which differs from the solid line for the parameter regime $\omega(\lambda,\rho_{a},\rho_{b})<-1/2$, which corresponds to $\lambda\lesssim-1.0$ in the figure.
Cumulants: The scgf (3) encapsulates all cumulants of $Q_{T}$ in the large $T$ limit for the semi-infinite SSEP. While the first three cumulants were initially reported in [2023_Saha_Current], we present here the fourth cumulant for $\rho_{a}=\rho_{b}=\rho$
$\frac{{\big<Q_{T}^{4}\big>}_{c}}{\sqrt{T}}\simeq\frac{4}{\sqrt{\pi}}\,\rho\,(1-\rho)\Big[1-12\,\big(\sqrt{2}-1\big)\,\rho\,(1-\rho)\Big].$ (26)
Infinite line SSEP with a slow bond: The scgf of current for an infinite line SSEP with a single slow bond can be obtained by treating the entire system as composed of three subsystems: two semi-infinite systems with unit hopping rates coupled to each other by a single slow bond with hopping rate $\Gamma/\sqrt{T}$. This leads to a variational formula [SM]
$\mu_{\text{inf}}^{\text{slow}}(\lambda,\rho_{a},\rho_{b})=$ $\max_{\rho_{0},\rho_{1}}\min_{\lambda_{0},\lambda_{1}}\Big[R_{\text{si}}\big(\omega(\lambda_{0},\rho_{a},\rho_{0})\big)$
$+\Gamma\,\omega(\lambda_{1}-\lambda_{0},$ $\rho_{0},\rho_{1})+R_{\text{si}}\big(\omega(\lambda-\lambda_{1},\rho_{1},\rho_{b})\big)\Big],$ (27)
where $\rho_{0(1)}$ is the density of the common site between the left (right) semi-infinite lattice and the slow bond. The analysis is similar to the discussion in [2021_Derrida_Large] for a related problem.
Figure 4: Scgf for infinite SSEP with a slow bond: The solid blue line denotes the scgf given in (24) for $\rho_{a(b)}=0.5$ and $\Gamma=1$. The blue markers denote the corresponding numerical results obtained by a continuous-time cloning algorithm with $T=500$ and $N_{c}=10^{4}$, for $\Gamma=1$. The solid red line denotes the $\Gamma\to\infty$ limit result [2009_Derrida_Current].
Figure 5: Scgf for semi-infinite SSEP with slow boundary coupling: The solid lines denote the scgf (23) for $\rho_{a(b)}=0.5$ and different values of $\Gamma$ (green for $\Gamma=1$, blue for $\Gamma=2$, and red for $\Gamma\to\infty$). The markers denote the corresponding numerical results obtained by the cloning algorithm with $T=500$ and $N_{c}=10^{4}$.
The expression (I) further simplifies using Eq. (29) of [2021_Derrida_Large] leading to the expression in (24). We confirm this theoretical result by comparing with numerical data from our rare event simulation as shown in Fig. 4.
A similar construction leads to the scgf (23) for semi-infinite SSEP with slow coupling to reservoir. A comparison of the result with numerical data is shown in Fig. 5.
A connection between SSEP and a kinetically constrained model: The energy-conserving spin-flip dynamics on a one-dimensional infinite lattice, defined in [1989_Spohn_Stretched], where each site $i$ at a given time $t$ is assigned with an Ising spin $S_{i}(t)=\pm 1$. The dynamics is such that a spin can flip only if its two nearest neighbor’s spins have opposite signs, therefore conserving the total energy (Fig. 6). This constraint ensures the conservation of number of domain walls, whose dynamics precisely mirror the particle movement in the SSEP on an infinite lattice.
This correspondence relates the spin-auto-correlation to the particle current in SSEP. We note that whenever a domain wall passes through a site $i$ (see Fig. 6), the associated Ising spin, $S_{i}$, changes its sign. If $Q_{t}$ is the net rightward flow of domain walls through the site $i$ in time $t$, then the quantity $S_{i}(0)S_{i}(t)$ takes values +1+1 or 1-1 for $Q_{t}$ being even or odd. This implies the auto-correlation
$\big<S_{i}(0)\,S_{i}(t)\big>=P_{t}(\text{even }Q_{t})-P_{t}(\text{odd }Q_{t}),$ (28)
with the probability $P_{t}(\text{even }Q_{t})$ of $Q_{t}$ being even and $P_{t}(\text{odd }Q_{t})$ for odd, respectively.
Figure 6: Dynamics of spins [1989_Spohn_Stretched] following a kinetic constraint of fixed energy: a spin can flip with unit rate only if its nearest neighbor’s spins are of opposite sign. For example, the fifth spin from left has left neighbor down and right neighbor up, so it is allowed to flip, making the domain wall move leftward by one lattice unit. The movement of domain walls (indicated by green lines) is analogous to the dynamics of particles in a SSEP.
The difference of the two probabilities in (28) relates to the generating function of current $\big<\mathrm{e}^{\lambda Q_{t}}\big$> in SSEP. Noting that $Q_{t}$ takes only integer values, we see that for $\lambda=\mathrm{i}\pi$,
$\big<\mathrm{e}^{\mathrm{i}\pi Q_{t}}\big>=\sum_{Q_{t}}P(Q_{t})\,\mathrm{e}^{\mathrm{i}\pi Q_{t}}=P_{t}(\text{even }Q_{t})-P_{t}(\text{odd }Q_{t}).$ (29)
From (28) and (29), it is evident that at large times, the spin-auto-correlation follows the asymptotics of the current statistics in SSEP. Using the results for the latter from [2009_Derrida_Current] we see that for large times $t$,
$\big<S_{i}(0)\,S_{i}(t)\big>\sim\mathrm{e}^{\sqrt{t}\mu_{\text{inf}}(\mathrm{i}\pi,\rho,\rho)}=\mathrm{e}^{-\sqrt{{t}/{\tau}}}$ (30)
with $\tau=[\mu_{\text{inf}}(\mathrm{i}\pi,\rho,\rho)]^{-2}$. Here, the density $\rho$ is the average uniform density of domain walls in the spin model [1989_Spohn_Stretched]. The exact result (30) is consistent with the bounds derived in [1989_Spohn_Stretched].